Solve these simultaneous equations using a algebric method
3x+7y=18
x+2y=5
x= y=

Found 2 solutions by sprolden, b_howie:
Answer by sprolden(40)

Solve these simultaneous equations using a algebric method
3x+7y=18
x+2y=5
x= y=
1- First, multiply the 2nd equation times -3
-3x-6y=-15
2- Add your equations
3x+7y = 18
-3x-6y =-15
3- Your solution should be y=3
4- Substitute 3 for y into either of the original equations
5- I'm going to use x+2y=5.
x+2(3)=5
x+6=5
x=-1
6- So x=-1 and y=3

Answer by b_howie(8)

3x+7y=18
x+2y=5

1.) solve one equation for x
x+2y=5
x=5-2y

2.) Plug in the value of x in the second equation
3x+7y=18
3(5-2y)+7y=5
15-6y+7y=5
1y=-10
y=-10

3.) Plug the value of y into either equation to solve for x
x+2y=5
x+2(-10)=5
x+(-20)=5
x-20=5
x=25

x=25
y=-10